OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)? @sasogeek @robtobey @Hero @Hero

OpenStudy (anonymous):

y2 - y1/x2 - x1 -4 - 6/-1 - 4 = -10/-5 = 2 So I take the 2 and plug it into y - 6 = 2(x - 4) y - 6 = 2x - 8 y = 2x - 2 So the slope is 2, since it's parallel?

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@reemii get ovaa here

OpenStudy (reemii):

yes. Just this is enough to find the slope: ""y2 - y1/x2 - x1 -4 - 6/-1 - 4 = -10/-5 = 2 """

OpenStudy (reemii):

Your equation at the beginning, when you don't have any information: y = mx + p. Now you know that m = -2. Your equation ,now: y = 2x + p. Use the fact the hte point (-1,1) is on the line to find the value of p.

OpenStudy (anonymous):

So wait, was I correct?

OpenStudy (reemii):

yes.

OpenStudy (anonymous):

What aboot this? I'm confused. http://prntscr.com/17g1dj

OpenStudy (reemii):

little area / big area

OpenStudy (anonymous):

x/x+5 = 4x/ x^2 + 7x + 10

OpenStudy (reemii):

little area: x(x+5). big area: (x^2+7x+10) * 4x there is a factorization that you can do with the polynomial x^2+7x+10.

OpenStudy (anonymous):

x(x+ 5)/(x + 5)(x + 2) = x/x + 2

OpenStudy (reemii):

yes. since there's also a "4x" factor, you get the second proposition.

OpenStudy (anonymous):

H-huh?